OpenStudy (anonymous):

−2(x + 3) = −2(x + 1) − 4

OpenStudy (anonymous):

-2x - 6 = -2x -2 - 4 -2x - 6 = -2x - 6 0 = 0 infinite amount of solutions

OpenStudy (anonymous):

Thank You soo much <3